OpenStudy (swaugh):

medal+ fan awaiting Marci finished the one hundred yard dash one and four-fifths seconds faster than Eva. If Eva finished in fourteen and seven-tenths seconds, how long did it take Marci to finish? eight and one-sixth seconds sixteen and one-half seconds twelve and nine-tenths seconds twenty-six and twenty-three fiftieths seconds

Directrix (directrix):

E - 1 4/5 = M where E is Eva's time and M is Marci's time 14 7/10 - 1 4/5 = M 14 7/10 - 1 8/10 = M

Directrix (directrix):

14 7/10 - 1 8/10 = M 13 17/10 - 1 8/10 = M 12 9/10 = M

Directrix (directrix):

@swaugh

OpenStudy (swaugh):

i got 16 1/2

Directrix (directrix):

Marcie finished faster than Eva so Marcie's time should be less than Eva's time of fourteen and seven-tenths.